OpenStudy (quickstudent):

Is this correct? question: How do you find the arc length? Explain the two methods learned in this section. my answer: There are two methods to find the length of an arc. 1. To find the length of an arc, or a part of a circle, if you are using a fraction such as 2/3 of the circle, you simply multiply the fraction times the circumference of the circle. 2. To find the length of an arc when you know the measure of the arc, such as 90 degrees, you would use the arc length formula: [(arc measure)/(360)] * C

OpenStudy (quickstudent):

@Awolflover1 @AloneS @phi @pooja195 @rishavraj @sleepyjess @vishal_kothari

OpenStudy (yanasidlinskiy):

Well are we supposed to explain it or we could just use of what we already know?.... Because... I know it this way...: \[s=r \theta \] s= arc length r=radius \[\frac{ x }{ 360 } =\frac{ y }{ 2\pi r }\] x= angle that spans arc length y= arc length

OpenStudy (quickstudent):

:o I didn't know there were other methods besides mine. That's why I asked so I could double check

OpenStudy (phi):

the s = r theta assumes theta is in radians